# Quantum DeepONet: Neural operators accelerated by quantum computing

### v1

@@ 4.2 Depolarizing noise on a RBS gate
The depolarizing noise is a widely adapted noise channel in analyzing the effects of quantum noise on variational quantum circuits [41, 58, 42, 59]. We take a closer look at how this type of noise affects a quantum layer composed of RBS gates and how the gate’s parameter values influence the level of error induced. Note that the statevector representation becomes insufficient to depict the quantum system under the influence of quantum noise. So, we use the density matrix to represent quantum states. A brief introduction to the definition and computation of the density matrix is provided in Appendix B.
@@
where $L=\{X,Y,Z,I\}$ is the set of Pauli matrices and the $2$-by-$2$ identity matrix $I$. In other words, the effect of $2$-qubit depolarizing noise means there is $161-15\lambda/$ 16 chance that the state $\rho$ remains unaffected, and an equal chance that any of the 15 possible $2$-qubit Pauli noise occurs on the state $\rho$. To further illustrate the impact of a noisy RBS gate, we consider a noise model where a noiseless RBS gate is first applied to the state $\rho$, followed by a depolarizing channel, resulting the final state $\rho^{\prime}$. In particular, we have
@@ 4.3 Noisy simulation results
In our subsequent investigation, we adapt two types of noise models to assess the accuracy of quantum layers under noisy conditions. The first approach, named simplified noise model, incorporates only 1-qubit and 2-qubit depolarizing noise channels on all basis gates. The goal of using this model is to examine the noise resilience of pyramidal circuits and the effects of the error mitigation technique on this well-researched noise channel, depolarizing noise channel. In the experiments, we select several different values for 1-qubit noise parameter $\lambda$, defined in Eq. (7), and set the 2-qubit noise parameter $\lambda^{\prime}:=0.8\lambda$. This is to guarantee both noise channels have the same error rate. Recall Eq. (7), 1-qubit depolarizing noise channel has the expression
@@
While the first approach aims to an direct and intuitive evaluation on the accuracy of pyramidal circuit under a noisy environment, depolarizing noise is insufficient to fully reflect the noise in real quantum computers and the 2-qubit gates usually have less fidelity than 1-qubit gates [62, 63, 64]. To fill this gap, we also carry experiments with the second approach: the backend-noise model from Qiskit Aer [54, 65]. The backend-noise model is in composite of
@@
The parameters of backend-noise model come from the regular benchmarking tests performed by the device vendor. By comparing these models, we can identify the feasibility of our quantum neural network and provide benchmarks for the improvement of near-term quantum computers.
@@ 4.3.1 Function approximation
We investigate the impact of finite-sampling error by varying the number of shots, $N_{shot}$, i.e., how many times we do the measurement to reconstruct the quantum state. The error with respect to true function value decreases when we increase number of shots (Fig. 6A). In this example, when the number of shots reaches $10^{8}$, shots-based simulation result is close to ideal simulation result. We further analysed the error between shots-based and ideal simulation (Fig. 6B), which is exactly the finite-sampling error mentioned in Section. 4.1. The error in Fig. 6B is proportional to $N^{-0.5}_{shot}$, which fits perfectly with the theoretical results in Eq. (6).
We further include depolarizing error, as introduced in Section.4.2, to estimate the affect of quantum gate noise. The error mitigation scheme is applied here. When $\lambda$ is within $[0,2\times 10^{-4}]$, the error increases almost linearly with $\lambda$ (Fig. 6C left). When we expand the range to $[0,2\times 10^{-3}]$, the error increases non-linearly and reaches a plateau at approximately $\lambda=10^{-3}$. As $\lambda$ reaches $2\times 10^{-3}$, the simulated error reaches around $20\%$ (Fig. 6C right).
Figure 6: Effect of quantum noise on function approximation example of $f(x)=1/(1+25x^{2})$. (A and B) Finite-sampling noise at different number of shots. (A) $L^{2}$ error between shots-based results and true function with different shots, compared with ideal simulation. (B) $L^{2}$ error between shots-based and ideal simulation. (C) Depolarizing noise model for different depolarizing parameter. In both cases, the number of shots is set to be $10^{7}$.
In order to simulate the performance of our quantum neural network on real quantum computer, we further adapt the backend-noise model in Qiskit. Here, we choose IBM_brisbane backend, loading the corresponding noise parameters for simulation. The error turns out to be $14.4\%$, suggesting some more sophisticated error mitigation methods will needed in future work. Since the scale of error is already too large in the simplest QOrthoNN experiment, the backend-noise model will not be tested in further experiments.
@@ 4.3.2 Antiderivative operator
Figure 7: Effect of noise on quantum DeepONet for the example of antidetivative operator. (A) $L^{2}$ relative error between shots-based and ideal simulation results for different layers. (B) Proportion of useful shots in total shots at different depolarizing noise level $\lambda$ when implementing error mitigation. (C and D) $L^{2}$ error between simulation results at different depolarizing level $\lambda$ and true solution. (C) Error mitigation is used. (D) Error mitigation is disabled. (E and F) The error for different neural network size. We set $\lambda=10^{-4}$ for all gates and fixed the number of shots at $10^{7}$. For each neural network sizes, we performed classical training 5 times until the test error is reduced to $3\%$. Each training run is quantumly simulated 3 times. The average and uncertainty of all 15 noisy simulation results were then calculated. (E) The network depth of both the branch and trunk nets is fixed at 5, while the width of both is varied simultaneously. (F) The network width of both the branch and trunk nets is fixed at 10, while the depths of both networks are varied simultaneously.
The impact of quantum noise on quantum DeepONet is also investigated using the antiderivative operator example (Section. 3.2) when $l=1.0$. In Fig. 7A, we plot the simulated finite-sampling error the scale of which is proportional to $N^{-0.5}_{shot}$, as expected in Eq. (6). When depolarizing quantum noise is considered, the error mitigation method discussed at Section. 2.1.4 can be applied. Although error mitigation helps eliminate undesired results caused by quantum noise, it also reduces the number of shots that are ultimately usable. It is obvious that
@@
The post-selection on the measurement outcomes significantly reduces the error in noise cases with both finite-sampling and depolarizing noise, comparing Fig. 7C with D. In Fig. 7C, where error mitigation is not applied and both unary and non-unary results are accepted, the error shows little reduction as number of shots increases. This is because the finite-sampling error is relatively insignificant under the influence of depolarizing error in this example, as can be seen by comparing the scales of errors in Figs. 7A and D. Therefore, increasing shots, which only reduce the finite-sampling error, does effective without error mitigation. However, with error mitigation, the overall error is significantly lower, making the reduction of finite-sampling error more obvious in the plot (Fig. 7D).
We also investigated how the network size can affect the error of noisy model (Figs. 7E and F). For each neural network size, we performed classical training 5 times. The networks were trained until the test error was reduced to $3\%$. For each trained network, we quantum-simulated 3 times. The parameters of simulations included $10^{7}$ shots and $\lambda=10^{-4}$ for depolarizing noise. It is important to note that even though the test error remained the same across classical training runs, the noisy simulation results varied. We believe this variation arises because the network converge to different parameter values in each training run, leading to different levels of error due to depolarizing noise. This observation aligns with our discussion in Section 4.2, which shows the parameters of RBS gates also influence the magnitude of errors. By comparing the two plots, we conclude that the error increases almost exponentially with increasing network depth. In contrast, when only the network width is increased, the error shows minimal growth within our experiment range. Therefore, quantum DeepONet shows resilience to noise with respect to network width. In practice, to minimize quantum noise, it is advisable to opt for wider rather than deeper neural networks.
@@ 5 Conclusions
We proposed quantum DeepONet, which can be both data-driven and physics-informed. Experimental results was conducted to confirm that quantum DeepONet performs efficiently in solving different PDEs. We further considered the impact of different noise model in simulation, and benchmarked the noise level and corresponding accuracy.
There are a few limitations in our current implementation. Based on the unary encoding, the quantum DeepONet currently could not handle large network width due to the limitation on the number of qubits and connectivities in the existing quantum devices, and the in-effectiveness of simulation on classical computers. However, although such demand on the number of qubits can be greatly reduced by removing the unary encoding, the increased cost of data loading and data tomography resulting from this change will require further analysis. On the other hand, in the noise simulation, both tested noise models do not include coherent noise and non-local noise such as cross-talk. A more complicated and realistic noise model are needed to examine the noise resilience of our design. Additionally, our future work will explore extending quantum DeepONet to accommodate more advanced architectures [66, 67], which will allow us to address a broader range of applications.

### v2

@@ 4.2 Depolarizing noise on a RBS gate
The depolarizing noise is a widely adapted noise channel in analyzing the effects of quantum noise on variational quantum circuits [41, 59, 42, 60]. We look closer at how this type of noise affects a quantum layer composed of RBS gates and how the gate’s parameter values influence the level of error induced. Note that the statevector representation becomes insufficient to depict the quantum system under the influence of quantum noise. So, we use the density matrix to represent quantum states. A brief introduction to the definition and computation of the density matrix is provided in Appendix B.
@@
where $L=\{X,Y,Z,I\}$ is the set of Pauli matrices and the $2$-by-$2$ identity matrix $I$. In other words, the effect of $2$-qubit depolarizing noise means there is $161-15\lambda/$ 16 chance that the state $\rho$ remains unaffected, and an equal chance that any of the 15 possible $2$-qubit Pauli noise occurs on the state $\rho$. To further illustrate the impact of a noisy RBS gate, we consider a noise model where a noiseless RBS gate is first applied to the state $\rho$, followed by a depolarizing channel, resulting in the final state $\rho^{\prime}$. In particular, we have
@@ 4.3 Noisy simulation results
In our subsequent investigation, we adopt two types of noise models to assess the accuracy of quantum layers under noisy conditions. The first approach, the simplified noise model, incorporates only 1-qubit and 2-qubit depolarizing noise channels on all basis gates. The goal of using this model is to examine the noise resilience of pyramidal circuits and the effects of the error mitigation technique on this well-researched noise channel, depolarizing noise channel. In the experiments, we select several different values for 1-qubit noise parameter $\lambda$, defined in Eq. (7), and set the 2-qubit noise parameter $\lambda^{\prime}:=0.8\lambda$. This is to guarantee both noise channels have the same error rate. Recall Eq. (7), 1-qubit depolarizing noise channel has the expression
@@
While the first approach aims for a direct and intuitive evaluation of the accuracy of pyramidal circuit under a noisy environment, depolarizing noise is insufficient to fully reflect the noise in real quantum computers and the 2-qubit gates usually have less fidelity than 1-qubit gates [63, 64, 65]. To fill this gap, we also carry out experiments with the second approach: the backend-noise model from Qiskit Aer [54, 66]. The backend-noise model is in composite of
@@
The parameters of the backend-noise model come from the regular benchmarking tests performed by the device vendor. By comparing these models, we can identify the feasibility of our quantum neural network and provide benchmarks for the improvement of near-term quantum computers.
@@ 4.3.1 Function approximation
We investigate the impact of finite-sampling error by varying the number of shots, $N_{shot}$, i.e., how many times we do the measurement to reconstruct the quantum state. The error with respect to true function value decreases when we increase the number of shots (Fig. 7A). In this example, when the number of shots reaches $10^{8}$, the shots-based simulation result is close to the ideal simulation result. We further analyzed the error between shots-based and ideal simulation (Fig. 7B), which is exactly the finite-sampling error mentioned in Section 4.1. The error in Fig. 7B is proportional to $N^{-0.5}_{shot}$, which fits perfectly with the theoretical results in Eq. (6).
We further include depolarizing error, as introduced in Section 4.2, to estimate the effect of quantum gate noise. The error mitigation scheme is applied here. When $\lambda$ is within $[0,2\times 10^{-4}]$, the error increases almost linearly with $\lambda$ (Fig. 7C left). When we expand the range to $[0,2\times 10^{-3}]$, the error increases non-linearly and reaches a plateau at approximately $\lambda=10^{-3}$. As $\lambda$ reaches $2\times 10^{-3}$, the simulated error reaches around $20\%$ (Fig. 7C right).
Figure 7: Effect of quantum noise on function approximation example of $f(x)=1/(1+25x^{2})$. (A and B) Finite-sampling noise at different number of shots. (A) $L^{2}$ error between shots-based results and true function with different shots, compared with ideal simulation. (B) $L^{2}$ error between shots-based and ideal simulation. (C) Depolarizing noise model for different depolarizing parameters. In both cases, the number of shots is set to be $10^{7}$.
In order to simulate the performance of our quantum neural network on real quantum computer, we further adapt the backend-noise model in Qiskit. Here, we choose ibm_brisbane backend, loading the corresponding noise parameters for the simulation. The error turns out to be $14.4\%$, suggesting some more sophisticated error mitigation methods will needed in future work. Since the scale of error is already too large in the simplest QOrthoNN experiment, the backend-noise model will not be tested in further experiments.
@@ 4.3.2 Antiderivative operator
The impact of quantum noise on quantum DeepONet is also investigated using the antiderivate operator example (Section 3.2) when $l=1.0$. In Fig. 8A, we plot the simulated finite-sampling error the scale of which is proportional to $N^{-0.5}_{shot}$, as expected in Eq. (6). When depolarizing quantum noise is considered, the error mitigation method is discussed in Section 2.1.4 can be applied. Although error mitigation helps eliminate undesired results caused by quantum noise, it also reduces the number of shots that are ultimately usable. It is obvious that
@@
The post-selection on the measurement outcomes significantly reduces the error in noise cases with both finite-sampling and depolarizing noise, comparing Fig. 8C with D. In Fig. 8D, where error mitigation is not applied and both unary and non-unary results are accepted, the error shows little reduction as number of shots increases. This is because the finite-sampling error is relatively insignificant under the influence of depolarizing error in this example, as can be seen by comparing the scales of errors in Figs. 8A and D. Therefore, increasing shots, which only reduce the finite-sampling error, is not effective without error mitigation. However, with error mitigation, the overall error is significantly lower, making the reduction of finite-sampling error more obvious in the plot (Fig. 8C).
We also investigated how the network size can affect the error of the noisy model (Figs. 8E and F). For each neural network size, we performed classical training 5 times. The networks were trained until the test error was reduced to $3\%$. For each trained network, we quantum-simulated 3 times. The parameters of simulations included $10^{7}$ shots and $\lambda=10^{-4}$ for depolarizing noise. It is important to note that even though the test error remained the same across classical training runs, the noisy simulation results varied. We believe this variation arises because the network converges to different parameter values in each training run, leading to different levels of error due to depolarizing noise. This observation aligns with our discussion in Section 4.2, which shows the parameters of RBS gates also influence the magnitude of errors. By comparing the two plots, we conclude that the error increases almost exponentially with increasing network depth. In contrast, when only the network width is increased, the error shows minimal growth within our experiment range. Therefore, quantum DeepONet shows resilience to noise with respect to network width. In practice, to minimize quantum noise, it is advisable to opt for wider rather than deeper neural networks.
Figure 8: Effect of noise on quantum DeepONet for the example of antiderivate operator. (A) $L^{2}$ relative error between shots-based and ideal simulation results for different layers. (B) Proportion of useful shots in total shots at different depolarizing noise levels $\lambda$ when implementing error mitigation. (C and D) $L^{2}$ error between simulation results at different depolarizing level $\lambda$ and true solution. (C) Error mitigation is disabled. (D) Error mitigation is used. (E and F) The error for different neural network size. We set $\lambda=10^{-4}$ for all gates and fixed the number of shots at $10^{7}$. For each neural network size, we performed classical training 5 times until the test error is reduced to $3\%$. Each training run is quantumly simulated 3 times. The average and uncertainty of all 15 noisy simulation results were then calculated. (E) The network depth of both the branch and trunk nets is fixed at 5, while the width of both is varied simultaneously. (F) The network width of both the branch and trunk nets is fixed at 10, while the depths of both networks are varied simultaneously.
@@ 5 Conclusions
We proposed quantum DeepONet, which can be both data-driven and physics-informed. Experimental results were conducted to confirm that quantum DeepONet performs efficiently in solving different PDEs. We further considered the impact of different noise models in simulation and benchmarked the noise level and corresponding accuracy.
There are a few limitations in our current implementation. Based on the unary encoding, the quantum DeepONet currently could not handle large network width due to the limitation on the number of qubits and connectivities in the existing quantum devices, and the in-effectiveness of simulation on classical computers. Also, the circuit in the quantum layer still has the $\mathcal{O}(n)$ depth, which is beyond the $\mathcal{O}(\log n$) bound on producing entanglement on NISQ devices [67]. On the other hand, although such demand on the number of qubits can be greatly reduced by removing the unary encoding, the increased cost of data loading and data tomography resulting from this change will require further analysis. It may require a design to apply non-linear transformation without collapsing the quantum state, such as the method in [68]. In terms of experiment settings in the noise simulation, both tested noise models do not include coherent noise and non-local noise such as cross-talk. A more complicated and realistic noise model is needed to examine the noise resilience of our design. Additionally, our future work will explore extending quantum DeepONet to accommodate more advanced architectures [69, 70], which will allow us to address a broader range of applications.